OpenStudy (safia21):

Write the standard form of the equation of the line described. through: (2,0) , parallel to y = 2/3x

jimthompson5910 (jim_thompson5910):

y = mx+b y = (2/3)x+b 0 = (2/3)(2) + b 0 = 4/3+b -4/3 = b b = -4/3 So the equation is \[\large y = \frac{2}{3}x-\frac{4}{3}\]